### v2

# The Role of Deductive and Inductive Reasoning in Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in reasoning tasks, yet their reliance on static prompt structures and limited adaptability to complex scenarios remains a significant challenge. In this paper, we propose the Deductive and InDuctive(DID) method, a novel framework that enhances LLM reasoning by dynamically integrating both deductive and inductive reasoning approaches. Drawing from cognitive science principles, DID implements a dual-metric complexity evaluation system that combines Littlestone dimension and information entropy to precisely assess task difficulty and guide decomposition strategies. DID enables the model to progressively adapt its reasoning pathways based on problem complexity, mirroring human cognitive processes. We evaluate DID’s effectiveness across multiple benchmarks, including the AIW and MR-GSM8K, as well as our custom Holiday Puzzle dataset for temporal reasoning. Our results demonstrate significant improvements in reasoning quality and solution accuracy - achieving 70.3% accuracy on AIW (compared to 62.2% for Tree of Thought), while maintaining lower computational costs. The success of DID in improving LLM performance while preserving computational efficiency suggests promising directions for developing more cognitively aligned and capable language models. Our work contributes a theoretically grounded, input-centric approach to enhancing LLM reasoning capabilities, offering an efficient alternative to traditional output-exploration methods.
The Role of Deductive and Inductive Reasoning in Large Language Models
Chengkun Cai1∗,  Xu Zhao1∗,  Haoliang Liu2∗,  Zhongyu Jiang3,  Tianfang Zhang4, Zongkai Wu5,  Jenq-Neng Hwang3,  Serge Belongie6,  Lei Li3,6† thanks: * Equal contribution. thanks: $\dagger$ Corresponding author: lilei@di.ku.dk 1University of Edinburgh  2University of Manchester  4Tsinghua University 5FancyTech  3University of Washington  6University of Copenhagen
1 Introduction
Large Language Models (LLMs), such as GPT-4, have transformed natural language processing by excelling in tasks such as language translation, summarization, and question-answering (OpenAI, 2023), particularly in reasoning tasks and few-shot learning. However, their reliability in problem-solving remains debatable. While Zhou et al. (2024) notes that scaling and fine-tuning can introduce unpredictable errors even in simple tasks, recent methodologies like Chain of Thought (CoT) (Wei et al., 2022b) have shown significant improvements in arithmetic and symbolic reasoning tasks (Li et al., 2024). Studies have demonstrated that LLMs can achieve high accuracy in multi-step reasoning when guided by structured approaches like CoT and self-consistency (Bubeck et al., 2023; Wang et al., 2022). Additionally, techniques such as reinforcement learning from human feedback (RLHF) have proven effective in reducing harmful or inaccurate outputs (Ouyang et al., 2022; Christiano et al., 2017).
Despite these advances, LLMs face significant challenges with complex and evolving tasks due to their reliance on static prompt structures and pre-learned patterns. This limitation manifests in tasks requiring logical reasoning, such as calculating family relationships or performing numerical comparisons (Nezhurina et al., 2024). Unlike human problem-solving, which dynamically adjusts strategies based on task complexity through inductive and deductive reasoning (Sloman, 2009), LLMs often struggle to adapt their reasoning processes to novel situations (Marcus, 2020; Hendrycks et al., 2020).
This adaptability gap becomes particularly evident in tasks requiring dynamic adjustment or incremental problem-solving. While existing approaches like CoT (Wei et al., 2022b), Tree-of-Thought (ToT) (Yao et al., 2024), Temperature-Tree-of-Thought (T2oT) (Cai et al., 2024), and Graph-of-Thought (GoT) (Besta et al., 2024) have made progress through extensive output exploration, they often incur significant computational costs. For instance, ToT achieves 62.2% accuracy on the AIW benchmark but requires substantial output token generation for exploring multiple reasoning paths, resulting in higher computational overhead (\$0.0038 per case compared to \$0.0022 for CoT).
Figure 1: Performance comparison of different reasoning approaches (IO, CoT, ToT, and DID) across model complexity. The x-axis represents an estimated measure of complexity that considers both model size (following public estimates from Abacha et al. (2024)) and reasoning token cost. The y-axis shows accuracy on the AIW reasoning benchmark. The relative positioning of models on the complexity axis is based on their approximate parameter counts and the computational overhead required for inference.
To address these challenges, we propose the De-In-Ductive (DID) method, a novel approach that enhances LLM reasoning by integrating both inductive and deductive reasoning processes within the prompt construction framework. Unlike previous methods that focus on expanding output exploration, DID takes an input-centric approach inspired by Test-Time Training techniques, strategically investing in input structuring to enable more efficient reasoning. The DID framework incorporates two key innovations: problem complexity evaluation and dynamic reasoning adjustment. For problem complexity evaluation, we introduce a dual-metric system that considers both the Littlestone dimension (measuring structural problem complexity) and information entropy (quantifying instance problem complexity) of problems, enabling precise assessment of task difficulty and guiding decomposition strategy.
Grounded in cognitive science models of human reasoning, DID implements a hybrid approach that mirrors human cognitive strategies. The method operates in two phases: first, it employs inductive reasoning to derive general rules from specific instances, progressively increasing problem complexity while maintaining similar Littlestone dimension; then, it applies deductive reasoning to solve particular problems, where the dynamic reasoning adjustment mechanism leverages problem complexity assessment to adaptively control the reasoning chain length and decomposition granularity.
We validate DID’s effectiveness on established benchmarks including AIW and MR-GSM8K (Wei et al., 2022a; Cobbe et al., 2021), as well as our custom Holiday Puzzle dataset focusing on holiday date calculations. As shown in Figure 1, our empirical results demonstrate significant improvements in both solution accuracy and reasoning quality, achieving 70.3% accuracy on AIW (compared to 62.2% for ToT) while maintaining lower computational costs (\$0.0031 vs \$0.0038 per case). This work makes the following key contributions:
  • We propose an innovative input-centric approach to LLM reasoning through the De-In-Ductive (DID) framework, which differs from existing output-exploration methods by strategically investing in input structuring. This approach fundamentally changes how we enhance LLM reasoning capabilities, offering a more efficient alternative to traditional methods.
  • We develop a theoretically grounded complexity evaluation system that combines Littlestone dimension and information entropy, enabling precise assessment of task difficulty and guiding the dynamic integration of inductive and deductive reasoning processes.
  • Through extensive empirical evaluations across diverse reasoning tasks, we demonstrate that DID not only achieves superior accuracy but also maintains lower computational costs through efficient input utilization, establishing a new direction for efficient LLM reasoning enhancement.
2 Related Works
Cognitive Science and Deductive-Inductive Reasoning
Deductive and inductive reasoning are essential in cognitive science, with deductive reasoning applying general principles to specific cases, and inductive reasoning generalizing from observations. Cognitive models view these approaches as complementary: inductive reasoning generates hypotheses, while deductive reasoning tests them (Wason, 1960). This combination enhances problem-solving, especially in uncertain domains where balancing exploration and validation is key (Johnson-Laird, 1983; Tversky and Kahneman, 1974). Well-structured problems typically favor deductive reasoning, whereas ill-structured problems benefit from inductive reasoning (Funke, 2013). Cognitive science insights have been integrated into neural networks, improving generalization (L Griffiths et al., 2008; Tenenbaum et al., 2011).
LLMs for Reasoning and Prompting Techniques
While LLMs like GPT-4 excel at tasks such as text generation, they struggle with logical reasoning and complex deduction (OpenAI, 2023; Nezhurina et al., 2024). Techniques like CoT (Wei et al., 2022b), ToT (Yao et al., 2024), and GoT (Besta et al., 2024) improve reasoning by structuring problems, but they require extensive prompt engineering and lack real-time adaptability. The DID framework addresses these limitations by dynamically integrating inductive and deductive reasoning, improving adaptability and consistency in complex tasks (Marcus, 2020; Gershman et al., 2015).
Recent advances in LLM reasoning capabilities have explored different approaches to enhancing model performance at test time. Models like DeepSeek-R1 Guo et al. (2025) and o1-ioi El-Kishky et al. (2025) achieve impressive results by leveraging reinforcement learning during pre-training and extending reasoning paths during inference - with DeepSeek-R1 reaching 79.8% accuracy on AIME 2024 through naturally emerged reasoning behaviors, and o1-ioi employing sophisticated test-time compute strategies to evaluate multiple solution candidates. Our DID framework takes a complementary approach by explicitly structuring input complexity during inference. This framework maintains explicit control through structured decomposition without requiring extensive computational resources or complex test-time evaluation strategies, potentially offering a more efficient path to enhanced reasoning capabilities while remaining easily accessible to the broader research community.
Inductive Inference and Online Learning
Recent work links inductive inference to online learning theory. (Lu, 2024) demonstrate that inductive inference is possible for hypothesis classes decomposable into countable unions with finite Littlestone dimension. This result extends classical induction models, such as Solomonoff’s (Solomonoff, 1964). The connection between Littlestone dimension and learning complexity informs the DID framework, suggesting that decomposing tasks into simpler components can enhance learning and generalization.
3 Methodology
3.1 Problem Formalization and Complexity Evaluation
Most reasoning tasks encountered by LLMs can be characterized as sequential learning problems with finite Littlestone dimension. According to recent theoretical work, a hypothesis class is learnable through inductive inference if and only if it can be decomposed into a countable union of classes with finite Littlestone dimension.
3.1.1 Littlestone Dimension and Beyond
For traditional online learning problems, the Littlestone dimension $d$ alone sufficiently characterizes problem difficulty. This dimension quantifies the intrinsic sequential learning complexity by measuring:
  • The minimal depth of decision trees needed for solving the problem
  • The number of key decision points in the reasoning process
However, when dealing with Large Language Models (LLMs), we observe that problems with identical Littlestone dimensions can exhibit significantly different difficulty levels. For example:
Example 1
Alice has 0 brothers and 1 sister. How many sisters does Alice’s brother have?
Example 2
Alice has 3 brothers and 6 sisters. How many sisters does Alice’s brother have?
Both problems share the same Littlestone dimension as they follow identical reasoning patterns. Yet, LLMs consistently perform better on Problem A. This discrepancy arises from several theoretical considerations:
  1. 1.
    Feature Vector Differences: In LLM’s internal representations, simpler numerical relationships create clearer, more distinguishable feature vectors
  2. 2.
    Distribution Shift: Larger numbers and more complex relationships often represent a shift from the training distribution
  3. 3.
    Information Bottleneck Theory: With increasing problem scale, the extraction of relevant information becomes more challenging due to the constrained capacity of intermediate representations
3.1.2 Information Entropy Component
To account for these LLM-specific challenges, we introduce an information entropy component $H$ that quantifies:
  • The complexity of numerical relationships
  • The density of relevant information that needs to be extracted
  • The scale of variables involved in the problem
For a problem instance $p$ with $n$ variables $\{x_{1},...,x_{n}\}$, we define its entropy as:
$H(p)=\log_{2}\left(\prod_{i=1}^{n}(1+|x_{i}|)\right)$ (1)
where $|x_{i}|$ represents the absolute value of each numerical variable in the problem. This formulation:
  • Grows logarithmically with problem scale
  • Remains bounded for reasonable problem sizes
  • Captures the intuition that larger numbers and more variables increase processing difficulty
3.1.3 Problem Complexity Evaluation
The overall complexity of a problem $p$ is then defined as:
$C(p)=d\cdot H(p)$ (2)
This combined measure allows us to:
  1. 1.
    Distinguish between problems of equal Littlestone dimension but different scale complexity
  2. 2.
    Better predict LLM performance on reasoning tasks
  3. 3.
    Guide the decomposition of complex problems into manageable subproblems
Dynamic Reasoning Based on Problem Complexity
Algorithm 1 Problem Decomposition in DID
1:
2:Problem $p$ with Littlestone dimension $d$
3:Problem complexity $C(p)=d\cdot H$
4:Step size parameter $a\in[1,C(p)]$ controlling decomposition granularity
5:
6:Sequence of subproblems with increasing complexity
7:function DecomposeProblem($p,d$)
8:     $N\leftarrow\left\lceil\frac{C(p)}{a}\right\rceil$
[…]
10:     $p_{\text{base}}\leftarrow\text{CreateBaseCase}(p,d-2)$
[…]
12:     for $i\leftarrow 1$ to $N$ do
13:         if $i<N/2$ then
14:              $d_{\text{current}}\leftarrow d-1$
15:         else
16:              $d_{\text{current}}\leftarrow d$
17:         end if
18:         $p_{\text{next}}\leftarrow\text{IncreaseCplx}(p_{\text{base}},d_{\text{current}})$
[…]
20:         $p_{\text{base}}\leftarrow p_{\text{next}}$
21:     end for
22:     return subproblems
23:end function
The Algorithm 1 formalizes our DID framework’s problem decomposition approach. At its core, the algorithm dynamically decomposes complex problems into a sequence of progressively challenging subproblems while managing both structural complexity (Littlestone dimension) and information density.
The decomposition process starts by creating a base case with reduced dimension ($d-2$), achieved by setting certain variables to zero. This simplification maintains the essential reasoning structure while reducing the problem’s complexity to its most basic form. From this foundation, the algorithm iteratively constructs $N=\lceil C(p)\rceil$ subproblems of increasing complexity, where $C(p)$ represents the overall problem complexity.
The algorithm employs a two-phase strategy in complexity progression:
  • In the first phase ($i<N/2$), it maintains a reduced dimension ($d-1$), allowing the model to establish fundamental patterns and relationships with minimal complexity
  • In the second phase ($i\geq N/2$), it restores the full dimension ($d$), gradually introducing complete problem complexity while building upon previously established patterns
This progressive approach mirrors human cognitive processes in problem-solving: starting with simplified versions, identifying core patterns, and systematically applying these insights to more complex cases. The IncreaseCplx function implements this gradual progression by introducing additional variables and relationships while maintaining the problem’s fundamental structure.
The algorithm’s dynamic dimension management ensures that the model can effectively balance between pattern recognition (inductive reasoning) in simpler cases and rigorous application (deductive reasoning) in more complex scenarios. This balance is crucial for maintaining both learning efficiency and solution accuracy across varying problem complexities.
3.2 De-In-Ductive (DID) Framework
Figure 2: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, Tree of Thought (ToT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
Figure 2 illustrates the comparison between the IO, CoT, and DID frameworks. The IO (Input-Output) Method processes natural language queries by retrieving patterns and facts without engaging in iterative reasoning. The Chain of Thought (CoT) Method improves logical reasoning by breaking down complex problems into sequential steps. Our proposed De-In-Ductive (DID) Method goes further by dynamically integrating inductive and deductive reasoning. By iteratively generating and testing hypotheses, DID adapts to problem complexities more effectively than static methods like CoT, optimizing problem-solving by balancing reasoning modes in response to task difficulty.
Dynamic Reasoning Based on Problem Complexity
Based on the problem complexity C(p), DID framework adaptively decomposes the problem and adjusts its reasoning process. For a typical problem with Littlestone dimension d (usually 3-5), we decompose it into subproblems:
  • Dimension Reduction: We maintain subproblems with dimension d or reduce to d-1 by fixing certain variables to 0, preserving the essential reasoning structure while reducing complexity
  • Progressive Complexity: Starting from simple cases with minimal information density, we gradually increase complexity by adding variables and relationships
  • Hierarchical Solution: Each subproblem ($K$) is solved using insights from solutions to the previous subproblem ($K-1$), enabling progressive knowledge accumulation
Inductive Reasoning
The inductive component enables pattern discovery and generalization from specific instances:
  • Pattern Recognition: Starting with simplified problem instances (reduced Littlestone dimension $d-2$ or $d-1$), the model identifies fundamental patterns and relationships. This aligns with the theoretical basis that inductive inference is possible when hypothesis classes have a finite Littlestone dimension.
  • Hypothesis Generation: Through progressive exposure to increasingly complex examples, the model generates and refines hypotheses about underlying problem structures. Each subproblem serves as a training instance for pattern recognition.
  • Complexity-Guided Learning: The inductive process is guided by the complexity measure $C(p)=d\cdot H$, ensuring that pattern recognition proceeds from simpler to more complex cases while maintaining manageable Littlestone dimensions.
Deductive Reasoning
The deductive component enables the systematic application of discovered patterns:
  • Rule Application: Once patterns are identified through induction, the model applies these rules deductively to solve more complex instances. This leverages the theoretical guarantee that hypothesis classes with finite Littlestone dimensions are learnable.
  • Verification Process: Each deductive step serves as a verification mechanism for inductively derived patterns, helping to refine and validate the model’s understanding.
  • Hierarchical Problem Solving: The deductive process follows the complexity hierarchy established during induction, ensuring that solutions build systematically on previously verified patterns.
The results from deductive applications inform and refine the inductive pattern recognition process, creating a continuous learning cycle that enhances the model’s problem-solving capabilities.
Integration with Existing Models
The De-In-Ductive (DID) method seamlessly integrates with various LLM architectures and existing techniques such as CoT prompting (Wei et al., 2022b). Through its structured framework combining inductive and deductive reasoning, DID enhances these methods by providing dynamic reasoning strategies and guided incremental reasoning, while maintaining computational efficiency. This approach creates a more flexible framework for LLMs to address complex problems without significant overhead.
4 Experiments
4.1 Experimental Setup
Baseline Methods and Models
We compare DID against three baseline prompting methods:
  • Input-Output (IO): directly utilizes the LLM without structured prompting
  • Chain of Thought (CoT): breaks down problems into sequential reasoning steps
  • Tree of Thought (ToT): explores multiple reasoning paths in a tree structure (T=3)
All methods are evaluated using three representative models:
  • GPT-4o and Claude 3.5 Sonnet: selected as two leading LLMs from different providers to demonstrate robustness across model architectures
  • GPT-3.5-turbo: included to evaluate method robustness across different model scales (in terms of parameter count)
For fair comparison, all model parameters (temperature, top-k sampling, etc.) are maintained at their default values. Evaluations are conducted in a zero-shot setting across all methods and models.
[…]
IO (%) CoT (%) ToT (%) DID (%) CoT (%) ToT (%) DID (%) IO (%) CoT (%) ToT (%) DID (%)
GPT-3.5 Turbo 6.7 8.6 7.2 13.3 68.1 74.0 73.3 0.2 1.4 2.0 5.6
GPT-4o 43.4 55.9 62.2 70.3 85.0 89.1 87.7 7.8 5.2 7.5 15.4
Claude 3.5 Sonnet 74.8 83.7 87.1 89.5 91.3 92.0 92.0 17.4 17.8 24.0 24.5
Table 1: Merged Results for GPT-3.5 Turbo, GPT-4o, and Claude 3.5 Sonnet across Different Tasks (Alice Problem, MR-GSM8K, Holiday Puzzle)
Table 2: GPT-4o Token Usage and Cost Comparison
Task Method Input/Output tokens Cost per case Accuracy (%)
Alice Problem IO 37/55 \$0.0007 43.4
CoT 45/210 \$0.0022 55.9
ToT 56/370 \$0.0038 62.2
DID 90/290 \$0.0031 70.3
MR-GSM8K CoT 86/1017 \$0.0104 85.0
ToT 91/1920 \$0.0194 89.1
DID 190/1230 \$0.0128 87.7
Holiday Puzzle IO 87/570 \$0.0059 7.8
CoT 96/1330 \$0.0135 5.2
ToT 110/2590 \$0.0262 7.5
DID 260/1740 \$0.0181 15.4
4.2 Tasks and Results
Alice Problems
Figure 3: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
The AIW dataset focuses on evaluating logical reasoning and deduction abilities through family relationship problems. We manually curated 113 unique problems after removing duplicates and existing prompts, with results averaging over 20 runs. In this task, DID demonstrates consistent superiority across all models:
  • GPT-3.5 Turbo: DID (13.3%) significantly outperforms IO (6.7%), CoT (8.6%), and ToT (7.2%)
  • GPT-4o: DID achieves 70.3% accuracy, surpassing IO (43.4%), CoT (55.9%), and ToT (62.2%)
  • Claude 3.5 Sonnet: DID reaches 89.5%, extending the lead over IO (74.8%), CoT (83.7%), and ToT (87.1%)
As illustrated in Figure 3, DID progressively guides LLMs through increasingly complex reasoning steps for family relationship problems. While traditional methods often fail by attempting to solve complex problems directly, DID breaks down the reasoning process into simpler subproblems, helping the model maintain logical consistency and avoid common errors. This structured approach enables LLMs to effectively handle complex relationship inference tasks.
As shown in Table 2, while DID requires slightly more input tokens (90 vs 56 for ToT), it maintains lower total computational costs (\$0.0031 vs \$0.0038) through more efficient output generation. This demonstrates the effectiveness of our input-centric approach in balancing performance and efficiency.
MR-GSM8K Math Problems
MR-GSM8K extends the GSM8K benchmark with meta-reasoning tasks, requiring models to identify and explain errors in provided solutions. Results show consistent performance across models:
  • GPT-3.5 Turbo: DID (73.3%) maintains competitive performance against CoT (68.1%) and ToT (74.0%)
  • GPT-4o: DID (87.7%) performs comparably to CoT (85.0%) and ToT (89.1%)
  • Claude 3.5 Sonnet: DID (92.0%) matches the strong performance of CoT (91.3%) and ToT (92.0%)
Figure 4: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
As shown in Table 2, DID achieves this performance with lower computational overhead than ToT (\$0.0128 vs \$0.0194), despite using more input tokens (190 vs 91). This efficiency gain comes from reduced output exploration needs.
Holiday Puzzle
This custom dataset comprises 20 holiday arrangement problems, testing models’ ability to calculate actual holiday days while accounting for weekends and compensatory workdays. Results demonstrate:
  • GPT-3.5 Turbo: DID (5.6%) outperforms IO (0.2%), CoT (1.4%), and ToT (2.0%)
  • GPT-4o: DID shows marked improvement (15.4%) over IO (7.8%), CoT (5.2%), and ToT (7.5%)
  • Claude 3.5 Sonnet: DID (24.5%) maintains advantage over IO (17.4%), CoT (17.8%), and ToT (24.0%)
The key to success in this task lies in discovering and applying the fundamental relationship Holiday rest days = Total rest days - Weekend rest days. As shown in Figure 4, baseline methods struggle with this pattern.
As shown in Table 2, while DID uses more input tokens (260 vs 110 for ToT), its efficient output generation results in lower total costs (\$0.0181 vs \$0.0262), demonstrating the scalability of our input-centric approach even in complex temporal reasoning tasks.
5 Conclusion
In this work, we introduced the De-In-Ductive (DID) method, a novel framework that dynamically integrates inductive and deductive reasoning to enhance the adaptability and reasoning capabilities of Large Language Models (LLMs). By leveraging cognitive science principles, the DID framework allows LLMs to evolve their problem-solving strategies in response to task complexity, overcoming the rigidity of static prompt structures. Through extensive empirical validation on both standard benchmarks and our custom Holiday Puzzle dataset, we demonstrated significant improvements in accuracy and reasoning quality, achieved without excessive computational costs. The success of DID in improving LLM reasoning while maintaining computational efficiency suggests promising directions for future research in making language models more cognitively aligned and capable of sophisticated reasoning.
6 Limitations
Despite the advances demonstrated by the DID framework, several important limitations and challenges remain to be addressed:
Fundamental Architecture Constraints
A key limitation lies in the fundamental architecture of LLMs. These models, based on next-token prediction, struggle to maintain coherent internal representations across multiple reasoning steps. While attention mechanisms allow reference to previous tokens, they lack robust cognitive structures for ensuring logical integrity throughout the reasoning process. This often leads to unexpected errors even in seemingly straightforward tasks.
Generalization Challenges
While DID shows strong performance on our evaluated tasks, ensuring consistent generalization to completely unseen problems remains challenging. The framework’s effectiveness may vary depending on the nature and complexity of new tasks, particularly those requiring novel forms of reasoning not encountered during development.